### v1

@@ 3.2 Script-Based Metrics
Given that each code cell in Jupyter notebooks is a regular Python script, many of the metrics for code scripts proposed by prior studies buse2009learning ; dorn2012general ; kasto2013measuring ; scalabrino2019automatically ; mathias1999role ; parker2003measuring are also applicable to code cells in Jupyter notebooks. Considering that a small percentage of notebooks use the concept of class and object orientation [^9] Based on the statistics of our notebooks dataset, only 3.5% of notebooks defined classes to implement objects. and that notebooks typically have a weaker structure, some metrics were ignored.
@@
  • Basic metrics: Basic metrics that are usually considered in most related articles are: the number of lines of code, number of blank lines of code, number of comment lines of code, number of comment words, number of statements, number of parameters and number of user-defined function buse2009learning ; dorn2012general ; kasto2013measuring ; scalabrino2019automatically .
  • Halstead metrics: Halstead metrics were employed by Parker and Becker parker2003measuring for the first time, to measure and compare the effectiveness of students’ solutions of two different code writing assessments based on the premise that the metrics can be seen as a measure of work done. Kasto kasto2013measuring introduced 4 of Halstead’s metrics as imperative metrics, which are: the number of operands, number of operators, number of unique operands and number of unique operators.
  • Cyclomatic Complexity: Cyclomatic complexity assigns a number for code complexity based on the code graph. It was first used by Mathias mathias1999role in order to examine the underlying nature of code designed to study the process of program comprehension.
  • KLCID: Kind of Line of Code Identifier Density (KLCID) klemola2003cognitive is one of the complexity metrics that analyzes the cognitive complexity of program comprehension tasks. The KLCID is calculated by counting the lines of conceptually unique code and calculating the density of identifiers. The KLCID computation involves transforming code into a format where lines with similar structures are identified. Unique lines with identifiers are counted, reflecting distinct patterns. KLCID is obtained by dividing the number of identifiers (such as variables, function names, or constants) by total lines, aiming to assess cognitive complexity. It recognizes that lines with more identifiers are harder to understand, emphasizing comprehension challenges posed by dense identifier usage in code segments.
4 Problem Definition
@@
In numerous studies sykes1983effect ; buse2009learning ; medeiros2018investigating ; scalabrino2019automatically , expert opinion has been employed as a criterion to measure CU. Questionnaires have been utilized as tools to assess the comprehension of experts based on their opinions regarding specific code snippets. In our study, we have also incorporated the criterion of expert opinion. However, our approach is unique as we analyze comments and opinions shared by individuals in notebook repositories. Nevertheless, two vital issues need to be addressed regarding these comments. Firstly, we must determine which comments are directly relevant to CU. Secondly, we need to assess the extent to which these comments have received attention from other experts. Taking these aspects into account, it is essential to define a criterion that effectively measures the ability to understand notebooks. Evaluating whether this criterion can independently assess CU or complement the previously established criteria is important. By comparing the results obtained from this criterion with a baseline, we can determine whether the new approach represents an improvement over existing methods, establishing its effectiveness and reliability.
RQ2: Is it feasible to develop a model capable of achieving satisfactory performance in assessing the code understandability of notebooks?
@@
Previous studies, such as scalabrino2019automatically ; buse2009learning have focused on developing methods for automatically measuring CU. However, these studies either did not provide a specific model or their model was limited to Java-based programs. In contrast, the primary objective of this study is to develop a model specifically for notebooks that can accurately estimate CU. In order to develop this model, two factors must be taken into consideration: the code metrics of the relevant notebooks for CU, and a suitable criterion for assessing CU based on the findings from RQ1. Previous studies have utilized upvotes as a metric to gauge the CU of large-scale notebooks. Therefore, a model trained using this criterion can serve as a valuable benchmark for comparison with our own model. If our method demonstrably outperforms previous methods on evaluation parameters, it can serve as an initial step toward addressing RQ3 in a positive manner.
RQ3: Which notebook metrics can affect CU, and to what extent can each metric influence code understandability?
After identifying the metrics related to CU and learning how to extract them (as described in Section 3), answering this research question requires addressing the following questions: How can we measure the impact of these metrics on notebook CU? Which metrics are the most effective in improving its understandability? Which group of metrics prove to be the most impactful in improving CU of notebooks? By addressing these research questions, we aim to pinpoint script-based and notebook-based metrics that could either enhance or hinder CU. By pinpointing the most influential metrics or groups of metrics, we can glean valuable insights on how to craft more comprehensible notebooks, thus fostering improved programming practices.
@@ 5 The Proposed Approach
Figure 1: Proposed Approach Overview to Predict Two Class of Notebooks: Normal Code Understandability (NCU) and Good Code Understandability (GCU).
@@
In fact, in Section 5.4, machine learning algorithms are used to find answers to RQ2 and RQ3. If we consider the learning operation as y=f(X), the output of Section 5.3 is used as X, and the output of Section 5.2 is used as y.
The machine learning algorithms in Section 5.4 utilize the findings from Section 5.3 as the input (X) and the results from Section 5.2 as the desired output (y). These algorithms aim is to find a function f (y=f(x)) that maps the input data X to the desired output y. This function can then be used to predict new, unseen data. Utilizing this learning operation, the algorithms aim to discover patterns or relationships between the notebooks’ code metrics and their CU, which can help provide answers to RQ2 and RQ3.
Figure 1 shows our approach’s steps to answer the research questions, which we will discuss in detail about each step.
@@ 5.1 Identifying Comments Related to CU
The comments corresponding to each notebook in the DistilKaggle dataset quaranta2021kgtorrent are available using Kaggle metadata [^10] https://www.kaggle.com/datasets/kaggle/meta-kaggle. The large number of comments available in our dataset gave us the opportunity to conduct an exploratory study and analyze these comments with the aim of establishing a metric for Jupyter notebooks’ CU.
During the data cleaning stage, any non-English comments were removed. We encountered a decline in performance during the training of our deep learning models, primarily attributed to the presence of lengthy comments within our dataset. Consequently, by filtering out comments longer than the 99th percentile of the dataset’s length distribution, we significantly reduced our training time while retaining the majority of our data. These comments had between 1,000 and 17,000 characters, which is less than 1% of the total data records.
@@
In the following, we outline our methodology for classifying user comments relating to Jupyter notebooks on Kaggle into three categories: “positive in terms of understandability,” “negative in terms of understandability,” or “no relation to understandability,” utilizing a team of annotators. Initially, we crafted a guideline outlining definitions for each category supplemented with illustrative examples. This guideline served to aid annotators in comprehending the process more effectively, thereby reducing discrepancies. We had a total of 659,894 comments. Based on calculations similar to those outlined by Zhu et al. zhu2021restoring , we determined that a sample size of 1,039 comments was required to attain a confidence level of 99% with a margin of error of ±4% sample2024 . Consequently, about 1,000 comments were extracted to train our model.
Four authors of this paper, having expertise in CU concepts, were selected as annotators for the labeling procedure. Each comment underwent assessment by two annotators who conducted an in-depth manual analysis of the content and assigned appropriate labels. An overall consensus was attained on 92% of the labels, resulting in a Cohen’s kappa coefficient of 80%, signifying reliable outcomes cohen1960coefficient . All contentious comments were meticulously reviewed, and a collaborative session was conducted with annotators to finalize the labeling decisions. Annotators engaged in productive discussions to harmonize and unanimously agree on the definitive labels.
@@
The distribution of these three groups of comments is shown in Figure 2. For a better understanding of how to label comments, several examples of comments and their labels are given in Appendix A.
@@
Therefore, out of the total 994 results, 467 were categorized as “related to code understandability” and 527 as “unrelated to code understandability”. To pinpoint comments specifically relevant to CU within the entire dataset, we fine-tuned the DistilBERT  [^11] DistilBERT is a smaller general-purpose language representation model and a distilled version of BERT, which is smaller, faster, computationally cheaper, and lighter. It can be used for text classification, named entity recognition, question answering, text generation, machine translation, and recommender systems. model on this data. DistilBERT was chosen due to its ability to effectively capture the contextual semantics of full sentences, which is crucial for understanding user comments. Additionally, as a distilled version of BERT, DistilBERT offers a good balance between performance and efficiency, making it suitable for our task. It is pre-trained on a large corpus of text, allowing it to understand the context of a sentence, and can be fine-tuned on a specific task, like classifying comments in this case. We had two main objectives in training this model. Firstly, we wanted to filter out comments that were unrelated to CU in order to evaluate the notebooks based only on their relevant comments. Secondly, we aimed to identify and include all comments related to CU when assigning an understandability score to the notebooks. The model achieved an F1-score of 90% and  accuracy of 92%, identifying 410,484 out of 659,894 comments in the whole dataset that were related to CU. Finally, a score of one was assigned to comments related to CU, while a score of zero was assigned to other comments.
@@ 5.2.1 Calculating CU Scores
As mentioned, the UOCU metric is utilized to measuring CU based on users’ opinions. To calculate UOCU for each notebook, we use Equation (3) which involves a function f that takes some metadata of a notebook and the BERTScore calculated in Section 5.1. This function then returns a new score called the “comment-based score”.
@@
- BERTScore is the binary score calculated for a notebook with DistilBERT,
@@
  • Apart from the relevance of a comment to the CU, the votes given to that comment show users’ agreement. To augment this impact, we multiply it by “(CommentVote + 1)”. Each comment starts with a value of one (+1), and with each additional upvote, this value increments by one.
  • We calculated the total scores ($\Sigma$) of comments for each notebook, which yielded the initial notebook score.
  • We divided the final notebook scores by the square root of the number of times the notebook was viewed on Kaggle. This was done to mitigate the impact of views on notebooks. This adjustment ensures that notebooks with higher view counts (and consequently more comments) do not receive disproportionately large scores, while notebooks with fewer view counts (and potentially fewer comments) do not receive unfairly lower scores.
  • We have introduced a “+$\alpha$ numerator” to differentiate notebooks with varying view counts that may not have CU-related comments. As illustrated in Equation (3), excluding the “+$\alpha$” would result in a numerator of zero for notebooks with zero comments, rendering them indistinguishable. With the inclusion of “+$\alpha$” in the numerator, we generate a small value that is then divided by the square root of the total views, creating diverse numbers. Through iterative experimentation, we have consistently observed that a value of 5 yields superior results in accomplishing this objective.
@@ 5.2.2 Converting CU Scores to Binary Class Labels
Up to this point, we have effectively formulated the UOCU benchmark leveraging DistilBERT. Should this criterion prove to be more fitting for comprehending notebook code compared to previous criteria(which is proved in the Section 6), it can be adopted as the label for the dataset, enabling the training of a model to assess the understandability of notebooks.
@@
In the following two paragraphs, the reason behind the converting the CU score to discrete class labels is elaborated. Linear regression is perhaps the most commonly used tool in statistics and as such is familiar to a diverse range of students and researchers. However, an even wider segment of the educated public do not understand regression or least squares and thus have to take many statistical results on faith, for example accepting results that are labeled as statistically significant without being able to interpret their numerical values. As a result, we would like to approximate the regression using a simple comparison. For this purpose, we can simplify the interpretation of regression by discretizing. In common practice, variables are discretized into two categories (i.e., the predictor x falling above or below some threshold) Andrew2009Splitting . However, as shown in this study Andrew2009Splitting , we can do better by discretizing the objective variable into three values and throwing away the middle category. The study shows the efficiency of this method. It makes the general recommendation that the high and low categories each be set to contain 1/4 to 1/3 of the data, which results in comparisons with approximately 80%–90% of the efficiency of linear regression.

### v2

@@ 3.2 Script-Based Metrics
  • Basic metrics: The number of lines of code, number of blank lines of code, number of comment lines of code, number of comment words, number of statements, number of parameters, and number of user-defined functions.
  • Halstead Metrics: Halstead metrics were first employed by Parker and Becker parker2003measuring to measure and compare the effectiveness of students’ solutions in two different coding assessments, based on the premise that these metrics can be viewed as indicators of work done. Kasto kasto2013measuring later identified four of Halstead’s metrics as imperative metrics: the number of operands (the total count of variables and constants in the code), the number of operators (the total count of actions performed by the code, such as arithmetic operations and control flow statements), the number of unique operands (the count of distinct variables and constants), and the number of unique operators (the count of distinct actions represented in the code).
  • Cyclomatic Complexity: Cyclomatic Complexity assigns a numerical value to code complexity based on the code’s control flow graph. It was first utilized by Mathias mathias1999role to investigate the underlying nature of code designed for studying the process of program comprehension.
  • KLCID: The Kind of Line of Code Identifier Density (KLCID) klemola2003cognitive is a complexity metric that analyzes the cognitive complexity associated with program comprehension tasks. The KLCID is calculated by counting the lines of conceptually unique code and determining the density of identifiers. The computation of KLCID involves transforming code into a format that identifies lines with similar structures. Unique lines containing identifiers are counted, reflecting distinct patterns in the code. KLCID is obtained by dividing the number of identifiers (such as variables, function names, or constants) by the total number of lines, thereby assessing cognitive complexity. This metric recognizes that lines with a higher density of identifiers are more challenging to understand, highlighting the comprehension difficulties posed by the use of dense identifiers in code segments.
4 Challenges and Research Directions
In the context of modern software development, computational notebooks have emerged as a vital tool for data scientists and researchers, particularly those using the Python programming language de2024bug . However, the quality of code and documentation within these notebooks often varies significantly, leading to challenges in code understandability and maintainability. This study seeks to assess code quality in Jupyter Notebooks by detecting deficiencies in understandability using static analysis metrics and transformer-based language models, with static metrics serving as inputs to our model. Furthermore, we provide insights into which factors, as captured by our metrics, most significantly influence CU.
@@ 4 Challenges and Research Directions
In numerous studies sykes1983effect ; buse2009learning ; medeiros2018investigating ; scalabrino2019automatically ; scalabrino2019automatically , expert opinion has been employed as a criterion to measure CU. Questionnaires have been utilized as tools to assess the comprehension of experts based on their opinions regarding specific code snippets. Unlike traditional methods that primarily focus on code structure or static analysis, our study integrates user opinions as a core component in measuring code understandability. What distinguishes our approach is that we analyze comments and feedback directly shared by users within notebook repositories, providing an additional layer of insight. This inclusion captures a broad spectrum of real-world perceptions and interactions from a large user base, which is often absent in methods that depend solely on structural metrics or small-scale, survey-based feedback. However, two critical challenges must be addressed regarding these comments. First, we need to identify which comments are directly relevant to CU. Second, we must assess the extent to which these comments have been acknowledged or validated by other users. Considering these factors, it is essential to establish a criterion that effectively measures a notebook’s understandability. Evaluating whether this criterion can independently assess CU or serve as a complement to previously established criteria is crucial. By comparing the outcomes derived from this new criterion against a baseline, we can determine if this approach represents an improvement over existing methods, thus confirming its effectiveness and reliability.
RQ2: Can a model be developed that effectively classify code understandability of notebooks based on the evaluation metrics established in this study?
@@
Previous studies, such as scalabrino2019automatically ; buse2009learning ; lavazza2023empirical , have focused on developing methods for automatically measuring CU. While these studies primarily utilized Java-based programs due to their influence on the programming community, they did not provide a specific model applicable to other programming languages. In order to develop this model, two factors must be taken into consideration: the code metrics relevant to notebook CU, and a suitable criterion for assessing CU based on the findings from RQ1. Previous studies have utilized upvotes as a metric to gauge the CU of large-scale notebooks datasets.
Therefore, a model trained using this criterion can serve as a valuable benchmark for comparison against our own model. If our method demonstrably outperforms previous methods on evaluation parameters, it can serve as an initial step toward addressing RQ3 in a positive manner.
RQ3: Which notebook metrics are associated with code understandability, and how significantly do they influence the classification of code understandability?
After identifying the metrics related to CU and learning how to extract them (as described in Section 3), addressing this research question requires considering the following points:
  • How can we measure the impact of these metrics on notebook CU?
  • Which metrics are the most effective in classifying notebooks based on CU?
  • Which group of metrics has the greatest impact on improving CU classification?
By exploring these questions, we aim to identify script-based and notebook-based metrics that either enhance or hinder CU. Understanding the most influential individual and grouped metrics will provide valuable insights into crafting more comprehensible notebooks, ultimately promoting better programming practices.
@@ 5 The Proposed Approach
Figure 1: Proposed Approach Overview to classify Two Class of Notebooks: Normal Code Understandability (NCU) and Good Code Understandability (GCU).
@@
In fact, in Section 5.4, we apply machine learning algorithms to address RQ2 and RQ3. These algorithms aim to identify relationships between input data and desired outcomes. The input data comes from the findings in Section 5.3, while the desired outcomes are derived from the results in Section 5.2.
The purpose of the machine learning algorithms in Section 5.4 is to learn how to accurately translate input data from Section 5.3 into the corresponding outcomes from Section 5.2. Once this learning process is complete, the algorithms can be used to classify notebooks based on new, unseen data. By discovering patterns between the notebook code metrics and their CU, these algorithms help provide insights into RQ2 and RQ3.
Figure 1 outlines the four main steps of our approach to answer the research questions:
  • Identifying Comments Related to CU
  • Calculating Notebook Score
  • Preparing Dataset
  • Utilizing Machine Learning
To start, Identifying Comments Related to CU involves pinpointing comments that contribute to CU by reviewing the text annotations in code cells. This step filters out comments unrelated to comprehension, focusing specifically on those expressing positive or negative opinions about the code’s CU. These selected comments establish a baseline for assessing CU levels. Next, Calculating Notebook Score evaluates each notebook’s overall quality and clarity. Scores are assigned based on multiple factors, including comment frequency, relevance to CU, and the number of views for each Jupyter notebook. These scores provide a quantitative assessment of each notebook’s understandability, facilitating effective comparisons across notebooks. The final scores also serve as labels for training machine learning models in subsequent steps. These steps, along with Preparing Dataset and Utilizing Machine Learning, are discussed in detail in Sections 5.1-5.4.
@@ 5.1 Identifying Comments Related to CU
Numerous studies have leveraged user opinions for various software engineering activities. Lin et al. lin2022opinion , in their systematic review of opinion mining for software development, identified over 70 papers from leading research databases, like panichella2015can ; fu2013people ; chen2014ar ; kurtanovic2017mining , that used user feedback to enhance software quality processes. These processes include evaluating software quality through crowd-sourced opinions, understanding general user satisfaction, assessing satisfaction with specific product features, identifying issues and feature requests from developer discussions and issue reports, and extracting actionable insights from user feedback.
In our study, we utilize user comments to detect CU in Jupyter notebooks. While user feedback may not always be wholly sufficient or perfectly consistent, it significantly complements expert analysis by providing a more comprehensive perspective on software understandability and user experience. To achieve this, we used comments associated with each notebook in the DistilKaggle dataset mostafavi2024DistilKaggle , accessible through Kaggle metadata [^7] https://www.kaggle.com/datasets/kaggle/meta-kaggle. The large number of comments available in our dataset gave us the opportunity to conduct an exploratory study and analyze these comments with the aim of establishing a metric for Jupyter notebooks’ CU.
During the data cleaning stage, any non-English comments were removed. We encountered a decline in performance during the fine-tuning of our deep learning models, primarily attributed to the presence of lengthy comments within our dataset. Consequently, by filtering out comments longer than the 99th percentile of the dataset’s length distribution, we significantly reduced our fine-tuning time while retaining the majority of our data. These comments had between 1,000 and 17,000 characters, which is less than 1% of the total data records.
@@
In the following, we outline our methodology for classifying user comments relating to Jupyter notebooks on Kaggle into three categories: “positive in terms of understandability”, “negative in terms of understandability”, or “no relation to understandability”, utilizing a team of annotators. Initially, we crafted a guideline outlining definitions for each category supplemented with illustrative examples, as outlined in Appendix B. This guideline served to aid annotators in comprehending the process more effectively, thereby reducing discrepancies. We had a total of 659,894 comments. Based on calculations similar to those outlined by Zhu et al. zhu2021restoring , we determined that a sample size of 1,039 comments was required to attain a confidence level of 99% with a margin of error of ±4% sample2024 . Consequently, 1,039 comments were extracted to fine-tune our model.
The team were selected as annotators for the labeling procedure comprised four of the paper’s authors, each with a solid background in data science and experience working with Jupyter notebooks. These individuals were selected due to their familiarity with code comprehension challenges and their capacity to evaluate the clarity and readability of code effectively. The composition of the annotation team ensured a consistent and informed approach to labeling, leveraging their collective knowledge to identify comments relevant to CU accurately.
Each comment underwent assessment by two annotators who conducted an in-depth manual analysis of the content and assigned appropriate labels. An overall consensus was attained on 92% of the labels, resulting in a Cohen’s kappa coefficient of 80%, signifying reliable outcomes cohen1960coefficient . All contentious comments were meticulously reviewed, and a collaborative session was conducted with annotators to finalize the labeling decisions. Annotators engaged in productive discussions to harmonize and unanimously agree on the definitive labels.
@@
The distribution of these three groups of comments is shown in Figure 2. For a better understanding of how to label comments, several examples of comments and their labels are given in Appendix C.
@@
Therefore, out of the total 994 results, 467 were categorized as “related to code understandability” and 527 as “unrelated to code understandability”. To pinpoint comments specifically relevant to CU within the entire dataset, we fine-tuned the DistilBERT, a smaller general-purpose language representation model and a distilled version of BERT, which is smaller, faster, computationally cheaper, and lighter model on this data. DistilBERT was chosen due to its ability to effectively capture the contextual semantics of full sentences, which is crucial for understanding user comments. Additionally, as a distilled version of BERT, DistilBERT offers a good balance between performance and efficiency, making it suitable for our task. It is pre-trained on a large corpus of text, allowing it to understand the context of a sentence, and can be fine-tuned on a specific task, like classifying comments in this case. We had two main objectives in fine-tuning this model. Firstly, we wanted to filter out comments that were unrelated to CU in order to evaluate the notebooks based only on their relevant comments. Secondly, we aimed to identify and include all comments related to CU when assigning an understandability score to the notebooks. We partitioned our dataset of $\sim 1000$ comments into training, validation, and test sets using a split ratio of 70-20-10. The F1 score and accuracy were computed on the test split, ensuring an unbiased evaluation of the model’s performance. The model achieved an F1-score of 90% and  accuracy of 92%. We then applied this model to the entire dataset. The model identified 410,484 out of 659,894 comments in the whole dataset that were related to CU. Finally, a score of one was assigned to comments related to CU, while a score of zero was assigned to other comments.
@@ 5.2.1 Calculating CU Scores
As mentioned, the UOCU criterion is utilized to measure CU based on users’ opinions. To calculate UOCU for each notebook, we use Equation (1) which involves a function f that takes some metadata of a notebook and the BERTScore calculated in Section 5.1.
@@
- BERTScore is the binary score calculated for a notebook comment with DistilBERT,
@@
Apart from the relevance of a comment to the CU, the votes given to that comment show users’ agreement. To augment this impact, we multiply it by “(CommentVote + 1)”. Each comment starts with a value of one (+1), and with each additional upvote, this value increments by one. We calculated the total scores ($\Sigma$) of comments for each notebook, which yielded the initial notebook score. We divided the final notebook scores by the square root of the number of times the notebook was viewed on Kaggle. This was done to mitigate the impact of views on notebooks. This adjustment ensures that notebooks with higher view counts (and consequently more comments) do not receive disproportionately large scores, while notebooks with fewer view counts (and potentially fewer comments) do not receive unfairly lower scores. We have introduced a “+$\alpha$ numerator” to differentiate notebooks with varying view counts that may not have CU-related comments. As illustrated in Equation (1), excluding the “+$\alpha$” would result in a numerator of zero for notebooks with zero comments, rendering them indistinguishable. With the inclusion of “+$\alpha$” in the numerator, we generate a small value that is then divided by the square root of the total views, creating diverse numbers. Through iterative experimentation, we have consistently observed that a value of 5 yields superior results in accomplishing this objective.
@@ 5.2.2 Converting CU Scores to Binary Class Labels
Up to this point, we have effectively formulated the UOCU criterion leveraging DistilBERT. Should this criterion prove to be more fitting for comprehending notebook code compared to previous criteria (which is proved in the Section 6), it can be adopted as the label for the dataset, enabling the training of a model to assess the understandability of notebooks.
@@
Continuous CU scores are less interpretable and cannot be directly used for classification. To address this, we can simplify the interpretation by discretizing. In common practice, variables are discretized into two categories (i.e., the predictor x falling above or below some threshold) Andrew2009Splitting . However, as shown in the study of Gelman et al. Andrew2009Splitting , we can do better by discretizing the objective variable into three values and throwing away the middle category. The study shows the efficiency of this method. It makes the general recommendation that the high and low categories each be set to contain 1/4 to 1/3 of the data, which results in comparisons with approximately 80%–90% of the efficiency of linear regression.